OpenStudy (anonymous):

List three different ways to write 5^11 as the product of two powers. Explain why all three of your expressions are equal to 5^11

OpenStudy (abb0t):

\(5 \times 5 \times 5 \times 5\times 5 \times 5\times 5 \times 5\times 5 \times 5 \times 5\)

OpenStudy (falco276):

Wait, that's the answer?

OpenStudy (falco276):

@abb0t He said 3 different ways. That's one.

OpenStudy (nubeer):

one more can be 5^6 x 5^5

OpenStudy (nubeer):

and u can pick one more just be changing the power of 5.. but just make sure that sum of powers would become 11

OpenStudy (nubeer):

its the same as i wrote.